The emf induced in the armature of a shunt generator is 596 V. The armature


resistance is 0.1Ω


(a) Compute the terminal voltage when the armature current is 460A.


(b) The field resistance is 110Ω. Determine the field current, and the current and power delivered to the external circuit



1
Expert's answer
2021-12-13T10:34:37-0500

a)

"i=\frac{\varepsilon }{\frac{\varepsilon}I-r}=500~mA,"

b)

"I_c=\frac{\varepsilon }R=5.4~A,"

"P={I_c^2}R=3.2~kW."